# IEEEexample:BSTcontrol

### v1

@@ Abstract
The increasing density of modern DRAM has heightened its vulnerability to Rowhammer attacks, which induce bit flips by repeatedly accessing specific memory rows. This paper presents an analysis of bit flip patterns generated by advanced Rowhammer techniques[1, 2] that bypass existing hardware defenses. First, we investigate the phenomenon of adjacent bit flips—where two or more physically neighboring bits are corrupted simultaneously—and demonstrate they occur with significantly higher frequency than previously documented. We also show that if multiple bits flip within a byte, they are more likely to be adjacent than randomly distributed: for example, if 4 bits flip within a byte, there is an 87% chance that they are all adjacent. We also demonstrate that bit flips within a row will naturally cluster together—likely due to the underlying physics of the attack.
We then investigate two fault injection attacks enabled by multiple adjacent or nearby bit flips. First, we show how these correlated flips enable efficient cryptographic signature correction attacks, successfully recovering ECDSA private keys from OpenSSL implementations where single-bit approaches would be unfeasible. Second, we introduce a targeted attack against large language models by exploiting Rowhammer-induced corruptions in tokenizer dictionaries of GGUF model files. This attack effectively rewrites safety instructions in system prompts by swapping safety-critical tokens with benign alternatives, circumventing model guardrails while maintaining normal functionality in other contexts. Our experimental results across multiple DRAM configurations reveal that current memory protection schemes are inadequate against these sophisticated attack vectors, which can achieve their objectives with precise, minimal modifications rather than random corruption.
@@ 1 Introduction
Our research addresses this gap by systematically analyzing adjacent bit flip occurrences using TRRespass and BlackSmith techniques. We investigate the physical mechanisms that increase the likelihood of correlated flips and quantify their probability distributions, demonstrating that there are likely underlying physical effects that cause bit flips to be clustered at the byte level and the row level. Furthermore, we explore the security implications of adjacent and clustered bit flips in two critical domains: cryptographic implementations and machine learning systems.
@@ 2 Background
Rowhammer Attack Mechanics The Rowhammer vulnerability exploits the physical limitations of DRAM by repeatedly activating (hammering) specific memory rows to induce bit flips in adjacent rows. This occurs because each activation introduces electrical disturbance that marginally depletes charge from nearby cells. While individual activations cause minimal disturbance, repeated activations within the refresh interval (typically 64ms) can accumulate sufficient disturbance to flip bits in victim rows.
@@
BlackSmith [2] further refined these techniques by introducing non-uniform hammering patterns that vary in both timing and access sequences. Unlike previous approaches that used fixed-interval activations, BlackSmith employs frequency-based hammering that optimizes the refresh-to-activation ratio for maximum effectiveness. This technique exploits the specific refresh patterns and timing vulnerabilities in TRR implementations, demonstrating successful bit flips in 30 of 40 tested DDR4 modules, including those resistant to TRRespass.
@@ 3 Localized Bit Flips
3.1 Rate of Incidence of Adjacent Flips
Figure 1: Heatmap showing probability of N adjacent bit flips being observed if M adjacent bit flips are observed (after profiling 100MB on DDR4 memory with BlackSmith[2])
Figure 1 presents a heatmap showing that the probability of observing that given M adjacent bit flips have occurred within a single byte in a particular row, N bit flips occur within another byte in the same row. This experiment was run on DRAM # A7 (see Appendix A for details on device types) DRAM memory on an Intel Coffee Lake Refresh CPU. Each cell represents how likely it is that when M adjacent flips are detected, N adjacent flips also appear. For example, if two adjacent bit flips are observed (M=2), there is a 96.8% probability of also encountering a single bit flip (N=1) elsewhere in the same row. The data reveals that single-bit flips (N=1) are highly probable across all cases, whereas larger N are more rare. This shows rows that contain at least 1 adjacent bit flip are likely to have another byte somewhere else in the row with a single bit flip.
Figure 2 shows the absolute number of adjacent bit flips recorded after conducting 1,000 fault attempts using BlackSmith fuzzing. The vertical axis is presented on a logarithmic scale. The results show that single-bit flips were more frequent, occurring  174k times. Two adjacent bit flips were observed  3k times, three adjacent bit flips appeared only 62 times, and four adjacent bit flips occurred twice, demonstrating that adjacent bit flipping is a real phenomenon, and even 4 adjacent bit flips can be seen after minimal fuzzing.
@@ 3.1 Rate of Incidence of Adjacent Flips
We ran a non-uniform Rowhammer access pattern as explained in [2] to explore rates of coincidence.
@@ 3.2 Distribution of Bit Flips within a Row
Having profiled substantial regions of memory as described in Sec. 3.1, we hypothesized that the distribution of all bit flips within a row might not be random. We chose to analyze the data at the byte level (at least 1 flip occurred within a byte) for two reasons; to simplify the the statistical analysis and computer hardware is architecturally designed around bytes. We devised a statistical model for the null hypothesis as: the distribution of bit flips is fully random, and the location of one flip does not impact the chance of a flip in any other location, and compared our observations to the predictions of the model. We found that bit flips are not randomly distributed, but somewhat clustered: bit flips are more likely to appear closer to other bit flips.
Null Hypothesis Consider a 8192-byte or 65536-bit row of DRAM known to have $n$ bit flips. If the locations of the bit flips are independent of each other, we may model the bits in the row as a series of Bernoulli trials with two outcomes (flip or no flip), where we estimate the fixed probability of a bit flip to be $p=\frac{n}{65536}$. Thus the probability distribution of the distance in bits $d$ between a bit flip and the next nearest bit flip is the geometric distribution
$(1-p)^{d-1}p\ \textrm{for}\ d\in\mathbb{N}=\{1,2,3,\ldots\}$ (1)
with mean distance $\frac{1}{p}=\frac{65536}{n}$.
@@ 3.3 Adjacent Bit Flips within a Byte
The theoretical probability of observing $k$ bit flips in a byte is:
$P_{adj}(n,k)=\frac{A(n,k)}{C(n,k)}$ (2)
For the case of $k=2$, we have $C(8,2)=28$ total combinations. The number of adjacent combinations is simply the number of adjacent pairs possible in 8 bits, which is 7 (positions 0-1, 1-2, 2-3, 3-4, 4-5, 5-6, and 6-7). Therefore:
@@
3 (n=565) 65.8% 10.7%
4 (n=23) 87.0% 7.1%
Observed Results Table 1 presents our experimental findings, which defy the idea that bit flips are randomly distributed within a byte. Instead, we show that if multiple bits flip within a byte, they are more likely to be adjacent than randomly distributed. For example, if 3 bits flip in a byte, we would expect them to be adjacent only 10.7% of the time, but instead we see them adjacent 65.8% the time. The results are even more apparent with 4 bit flips, where if we see 4 bit flips there is an 87% chance they are adjacent, where if the bits were randomly distributed there would only be a 7.1% chance they would be adjacent. While we did not see 5 adjacent bit flips, we believe 4 is enough to have significant impact in this study.
@@ 4.1 ECDSA Fault Injection
Our analysis focuses on OpenSSL’s ECDSA implementation, where we identified several locations vulnerable to adjacent bit flips. When signing a message, OpenSSL computes the signature as a pair $(r,s)$ where:
$s=k^{-1}(z+r\cdot d_{A})\mod n$
Here, $k$ is the secret nonce, $z$ is the message hash, $d_{A}$ is the private key, and $n$ is the order of the elliptic curve group. The security of the signature depends critically on the protection of both $d_{A}$ and $k$.
@@
The injection of adjacent flips with Rowhammer can create the opportunity for the signature correction scheme to provide the necessary data for this powerful modified BDD with predicate algorithm.
@@ 4.3 Conclusion
This paper demonstrated that modern Rowhammer attacks produce localized effects that have been previously underexplored. First, we discovered that two or more adjacent bits in memory could flip together more often than expected. We also showed that bytes containing a flip tend to cluster together in the same row. We used this to break cryptographic systems, stealing ECDSA private keys from OpenSSL with fewer mistakes than older methods. The second attack targeted large language models by corrupting the tokenizer file. This allowed attackers to modify safety instructions in system prompts without affecting the model’s normal behavior.
@@ Appendix A Testing Different Rowhammer Tools
We tested both TRRespass [1] and BlackSmith [2] and compared bit flip adjacency across multiple different DRAMs. The results of this study can be seen in Table 4.
@@ Appendix B Test Setup
A3, A4 Corsair CMU64GX4M4C3200C16 64GB
A7 Corsair CMK32GX4M2B3200C16 32GB
A8 G.SKILL F4-3600C16D-16GVKC 8GB×$\times$2

### v2

@@ Abstract
The increasing density of modern DRAM has heightened its vulnerability to Rowhammer attacks, which induce bit flips by repeatedly accessing specific memory rows. This paper presents an analysis of bit flip patterns generated by advanced Rowhammer techniques[1, 2] that bypass existing hardware defenses. First, we investigate the phenomenon of adjacent bit flips—where two or more physically neighboring bits are corrupted simultaneously—and demonstrate they occur with significantly higher frequency than previously documented. We also show that if multiple bits flip within a byte, we can probabilistically model the likelihood of flipped bits appearing adjacently. We also demonstrate that bit flips within a row will naturally cluster together—likely due to the underlying physics of the attack.
We then investigate two fault injection attacks enabled by multiple adjacent or nearby bit flips. First, we show how these correlated flips enable efficient cryptographic signature correction attacks, demonstrating how such flips could enable ECDSA private key recovery from OpenSSL implementations where single-bit approaches would be unfeasible. Second, we introduce a targeted attack against large language models by exploiting Rowhammer-induced corruptions in tokenizer dictionaries of GGUF model files. This attack effectively rewrites safety instructions in system prompts by swapping safety-critical tokens with benign alternatives, circumventing model guardrails while maintaining normal functionality in other contexts. Our experimental results across multiple DRAM configurations reveal that current memory protection schemes are inadequate against these sophisticated attack vectors, which can achieve their objectives with precise, minimal modifications rather than random corruption.
@@ 1 Introduction
Our research addresses this gap by systematically analyzing adjacent bit flip occurrences using TRRespass and BlackSmith techniques. We investigate the physical mechanisms that increase the likelihood of correlated flips and quantify their probability distributions, demonstrating that there are likely underlying physical effects that cause bit flips to be clustered at the row level. Furthermore, we explore the security implications of adjacent and high frequency bit flips in two critical domains: cryptographic implementations and machine learning systems.
@@ 2 Background
Rowhammer Attack Mechanics The Rowhammer vulnerability exploits the physical limitations of DRAM by repeatedly activating (hammering) specific memory rows to induce bit flips in adjacent rows. This occurs because each activation introduces electrical disturbance that marginally depletes charge from nearby cells. While individual activations cause minimal disturbance, repeated activations within the refresh interval can accumulate sufficient disturbance to flip bits in victim rows.
@@
BlackSmith [2] further refined these techniques by introducing non-uniform hammering patterns that vary in both timing and access sequences. Unlike previous approaches that used fixed-interval activations, BlackSmith employs frequency-based hammering that optimizes the refresh-to-activation ratio for maximum effectiveness. This technique exploits the specific refresh patterns and timing vulnerabilities in TRR implementations, demonstrating successful bit flips in 40 of 40 tested DDR4 modules, including those resistant to TRRespass.
Half Double [13] studied Rowhammer in LPDDR4x systems with on-die ECC, where single-bit errors are automatically corrected. In these systems, only double (or more) bit flips within an ECC codeword manifest as observable errors, as the ECC silently corrects single-bit flips. While Half Double adapted page table exploits to handle these multi-bit errors, their work focused on systems where observing multiple flips is a requirement due to ECC. In contrast, our work analyzes DDR4 without on-die ECC, where we can observe all bit flips including single-bit errors. This allows us to study the fundamental phenomenon: that when multiple bits flip, they exhibit strong spatial correlation and cluster adjacently at rates far exceeding random chance. Further, we demonstrate that logically adjacent bit flips—those at consecutive bit positions like $i$ and $i+1$—create unique exploitation opportunities. Unlike scattered multi-bit errors that simply bypass ECC, adjacent flips produce predictable arithmetic relationships (e.g., $\Delta=\pm 3\cdot 2^{i}$) that can be mathematically exploited in cryptographic attacks, as we show with ECDSA key recovery where adjacent bit patterns directly reveal nonce bits.
@@ 3 Localized Bit Flips
We define adjacent bit flips as bit flips occurring at consecutive bit positions within the logical address space of a byte (e.g., bits at positions i and i+1). While we acknowledge that logical adjacency may not correspond to physical adjacency in DRAM due to data swizzling  [16], our analysis focuses on the software-visible effects that are relevant to exploitation.
Figure 1 shows the absolute number of adjacent bit flips recorded after profiling $\sim$100Mb of memory using BlackSmith fuzzing. The vertical axis is presented on a logarithmic scale. The results show that single-bit flips were more frequent, occurring  174k times. Two adjacent bit flips were observed  3k times, three adjacent bit flips appeared only 62 times, and four adjacent bit flips occurred twice, demonstrating that adjacent bit flipping is a real phenomenon, and even 4 adjacent bit flips can be seen after minimal fuzzing.
@@ 3.1 Distribution of Bit Flips within a Row
Having profiled substantial regions of memory, we hypothesized that the distribution of all bit flips within a row might not be random. We chose to analyze the data at the byte level (at least 1 flip occurred within a byte) for two reasons; to simplify the the statistical analysis and computer hardware is architecturally designed around bytes. We devised a statistical model for the null hypothesis as: the distribution of bit flips is fully random, and the location of one flip does not impact the chance of a flip in any other location, and compared our observations to the predictions of the model. We found that bit flips are not randomly distributed, but somewhat clustered: bit flips are more likely to appear closer to other bit flips.
Null Hypothesis Consider a 8192-byte or 65536-bit row of DRAM known to have $n$ bit flips. If the locations of the bit flips are independent of each other, we may model the bits in the row as a series of Bernoulli trials with two outcomes (flip or no flip), where we estimate the fixed probability of a bit flip to be $p=\frac{n}{65536}$. Thus the probability distribution of the distance in bits $d$ between a bit flip and the next nearest bit flip is the geometric distribution $(1-p)^{d-1}p\ ~{}~{}~{}\textrm{for}\ ~{}d\in\mathbb{N}=\{1,2,3,\ldots\}$ with mean distance $\frac{1}{p}=\frac{65536}{n}$.
@@ 3.2 Adjacent Bit Flips within a Byte
The theoretical probability of observing $k$ bit flips in a byte is: $P_{adj}(n,k)=\frac{A(n,k)}{C(n,k)}$ For the case of $k=2$, we have $C(8,2)=28$ total combinations. The number of adjacent combinations is simply the number of adjacent pairs possible in 8 bits, which is 7 (positions 0-1, 1-2, 2-3, 3-4, 4-5, 5-6, and 6-7). Therefore:
@@
3 (n=565) 10.6% 10.7%
4 (n=23) 8.7% 7.1%
Observed Results Table 1 presents our experimental findings, which demonstrate that we can probabilistically model the likelihood of flipped bits appearing adjacently within a byte assuming flipped bits are distributed randomly though the byte. For example, we can see that given 10k bytes where 2 bits flipped, $\sim$25of those flipped bits appeared adjacently, matching our probabilistic model. For 3 and 4 bit flips within a byte, both the theoretical and observed frequency of adjacency are roughly equivalent as well.
@@ 4.1 ECDSA Fault Injection
Our analysis focuses on OpenSSL’s ECDSA implementation, where we identified several locations vulnerable to adjacent bit flips. When signing a message, OpenSSL computes the signature as a pair $(r,s)$ where: $s=k^{-1}(z+r\cdot d_{A})\mod n$. Here, $k$ is the secret nonce, $z$ is the message hash, $d_{A}$ is the private key, and $n$ is the order of the elliptic curve group. The security of the signature depends critically on the protection of both $d_{A}$ and $k$.
@@
While we demonstrate the feasibility of this attack through theoretical analysis, we acknowledge that full empirical validation remains future work. However, we show in theory that the injection of adjacent flips with Rowhammer can create the opportunity for the signature correction scheme to provide the necessary data for this powerful modified BDD with predicate algorithm.
@@ 4.2 LLM Dictionary Faulting
Impact of Adjacent Bit Flips on Token Complexity While single bit flips can create simple character substitutions as shown in Figure 3, adjacent bit flips expand the attack surface by enabling more complex token transformations. Table 3 presents examples of token swaps discovered after profiling 100MB of memory on DRAM A3 using [1]. These adjacent bit flip patterns enable semantically meaningful swaps such as "firearm" to "forearm" and "junction" to "function"—transformations that require coordinated multi-bit changes and demonstrate how localized bit flips can create sophisticated vocabulary corruptions beyond simple single-flip substitutions. This expanded capability increases the attack surface, as hundreds of additional unique token swaps become feasible when adjacent bit flips are considered, providing attackers with greater flexibility in crafting targeted manipulations of model behavior.
Table 3: Examples of complex token swaps enabled by adjacent bit flips after profiling 100MB of memory on DRAM A3 using [1]
Original Token Swapped Token
firearm forearm
junction function
dry Try
loyd load
scared soared
enter *inter
@@ 4.3 Conclusion
This paper demonstrated that modern Rowhammer attacks produce localized effects that have been previously underexplored. First, we discovered that modern Rowhammer methods generate many adjacent bit flips which we can probabilistically model. We also showed that bytes containing a flip tend to cluster together in the same row. We used this to break cryptographic systems, stealing ECDSA private keys from OpenSSL with fewer mistakes than older methods. The second attack targeted large language models by corrupting the tokenizer file. This allows attackers to modify safety instructions in system prompts without affecting the model’s normal behavior.
@@ Appendix A Testing Different Rowhammer Tools
We tested both TRRespass [1] and BlackSmith [2] and compared bit flip adjacency across multiple different DRAMs. The results of this study can be seen in Table 5. Note that this study was to demonstrate that the bit adjacency affect was present on multiple DIMMs; $\sim$1000 fault attempts generally represents $\sim$20Mb of scanned memory which is why the number of faults is less than in Table 1.
@@ Appendix B Test Setup
A3, A4 Corsair CMU64GX4M4C3200C16 16GB
A7 Corsair CMK32GX4M2B3200C16 16GB
A8 G.SKILL F4-3600C16D-16GVKC 8GB
@@ Appendix C Lattice Attacks on the Hidden Number Problem (HNP)
Appendix D Disclaimer
Andrew Adiletta’s affiliation with The MITRE Corporation is provided for identification purposes only, and is not intended to convey or imply MITRE’s concurrence with, or support for, the positions, opinions, or viewpoints expressed by the author. All references are public domain.Approved for Public Release; Distribution Unlimited. Public Release Case Number 25-1773 ©2025 The MITRE Corporation. ALL RIGHTS RESERVED.